Circular No. 5952

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1994E IN ANONYMOUS GALAXY C. Pollas, Observatoire de la Côte d’Azur, reports his discovery of a

supernova (m_{pg} = 19–20) located at α = 11^{h}29^{m}20^{s}.83, δ = +55^{o}38′08′′.3 (equinox 1950.0), which is 0′′.3 west and 3′′.6 south of the center of a spiral galaxy of m_{pg} ∼ 18. A nearby star (m_{pg} ∼ 15.5) is at (end figures) 22^{s}.89, 37′29′′.2. SN 1994E was detected on two technical pan films taken on Mar. 5.04 and 6.03 UT, while nothing is visible on a pair of similar films taken in early 1992; both pairs of films reach stellar mag 22.

A. V. Filippenko, T. Matheson, A. Dey, and H. Spinrad, University of California at Berkeley, report that an uncalibrated CCD spectrum (range 380–890 nm, resolution 1 nm), obtained on Mar. 13 UT with the W. M. Keck 10-m telescope, reveals that the object is a type-Ia supernova, roughly 2 months past maximum brightness. A spectrum of the parent galaxy yields a redshift of 0.06.

COMET McNAUGHT-RUSSELL (1993v) R. Schulz, Max-Planck-Institut für Aeronomie; D. D. Wellnitz, Uni-

versity of Maryland; and P. V. Birch, Perth Observatory, report: “We have imaged this comet with the 0.61-m Lowell Telescope (+ University of Maryland CCD camera system) at Perth Observatory on the six nights Mar. 9–14. The comet clearly shows short-term variations in brightness (the largest of which reach a factor of 1.6) in CN and C_{2}, as well as in broad red (RG665); the light curves in all three bands show the same pattern.”

SUPERNOVA 1994D IN NGC 4526 CCD photometry (Cousins system) by P. Heraudeau, P. Prugniel, and

J. Taupenas, using the 1.2-m telescope at Observatoire de Haute Provence: Mar. 11.08 UT, V = 13.48, B −V = −0.05, V −R = +0.15, V −I = +0.04; 13.05, 12.80, −0.09, +0.27 +0.21; 14.04, 12.53, −0.04, +0.25, +0.16.

NOVA Sgr 1994 = LUMINOUS RED VARIABLE IN SAGITTARIUS

Corrigendum. On IAUC 5949, line 24, for common phase read common envelope phase

1994 March 15 Daniel W. E. Green